please help me if the problem: f(x)=x-4 and g(x)= x^2+7x-5...
6.f o g 7. g o f

Answer by tutorcecilia(2152)

6. f(x) = x-4
f o g =(____) - 4
=(x^2+7x-5) -4 [dump g(x) into f(x) and simplify]
f o g =(x^2+7x-9)
.
.
7. g(x)= x^2+7x-5
g o f = (______)^2 + 7(__) -5
=(x-4)^2+7(x-4)-5 [dump f(x) into g(x) and simplify]
=x^2-8x+16+7x-28-5
=x^2-x-17